# November 5, 2025

### v1

@@ 1 Introduction
In recent work [1] we proposed a symplectic structure for the phase space of an arbitrary Lagrangian field theory:
@@
The formula assumes that the theory has been expressed in terms of a cyclic $L_{\infty}$ algebra, where $\omega$ is the Batalin-Vilkovisky inner product, $Q_{\Phi}$ is the kinetic operator around a solution $\Phi$, and $\sigma$ is a grade zero operator called the sigmoid. In this paper, and two more to appear soon [2, 3], we give some applications of this formula to string field theory (SFT). In principle the formula applies to both open and closed string field theory. However, we will focus on open (bosonic) strings. Closed SFT is a theory of gravity, which makes extraction of observables (such as energy) more subtle. Some relevant work on diffeomorphism invariance and boundary terms in closed SFT can be found in [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. We leave applications to closed SFT for the future.
An important motivation for our work is the analysis of Cho, Mazel, and Yin [14] who use an older proposal for the symplectic structure due to Witten [15] to compute the energy of rolling tachyon solutions on unstable D-branes [16]. Their results match with boundary state calculations carried out by Kudrna [17]. The agreement, however, requires discarding certain infinities contained within Witten’s formula. In this paper we would like to give a more coherent analysis using the new formula (1.1). In the process, we confirm that (1.1) defines a symplectic structure on the phase space of open string field theory.
The divergences in Witten’s symplectic structure stem from the use of the open string midpoint to measure time. In our calculations, we will measure time with the open string center of mass. This is the time coordinate associated to the component fields of the Fock space expansion of the SFT action. With this definition of time, the symplectic form (1.1) is finite. A further benefit is that its evaluation does not rely on special properties of the Witten vertex that have no analogue in other SFTs. A challenge, however, is that Witten’s string field theory is highly nonlocal in the center of mass coordinate. The nature of the nonlocality is also strange. The interactions are in a sense “more than local” in time, as expressed by couplings which grow super-exponentially at high energies. We refer to this as transgressive locality. A consequence is that fields do not interact in a well-defined way as a function of time. Instead interactions must be expressed in momentum space. The lack of a clear picture of time evolution is worrisome given that we are trying to do Hamiltonian mechanics, but we will see that everything is working.
@@
Finally we consider open string field theory. We compute the symplectic structure for rolling tachyon solutions in Siegel gauge using the center of mass coordinate to measure time. We reduce the expression to first subleading order to a collection of correlation functions on the surface defining the off-shell 4-point amplitude in Siegel gauge. The correlation functions are evaluated using squeezed state oscillator methods developed by Taylor [21], showing good agreement with the results of Cho, Mazel and Yin and very close agreement with boundary state results of Kudrna. The details of our numerical implementation can be found in the Mathematica file accompanying this paper. We conclude by reviewing how the divergences of Witten’s symplectic structure can be eliminated by using the lightcone coordinate of the string midpoint to measure time [22, 23]. Independence from the choice of sigmoid can then be used to explain why our results agree with those of Cho, Mazel, and Yin.
@@ 2 Transgressive locality
The cubic term contains an infinite derivative operator $e^{\frac{\Lambda}{2}\Box}$ which appears (for some $\Lambda$) ubiquitously in interactions between any fields in the Fock space expansion of any covariant string field theory. For this reason nonlocality in string field theory takes a fairly universal form. The infinite derivative operator can be interpreted as a formal power series in $\Box$. But a better definition, which is nonperturbative in $\alpha^{\prime}$, can be given in one of two ways: as a position space integral transform,
@@
The problem with momentum space is that we cannot talk directly about time evolution. This creates some worry, as the consistency of the symplectic structure relies on the assumption that fields propagate and interact in a somewhat localized fashion as a function of time. It might appear that SFT interactions are not like this. However, we will argue that really they are. In fact, in a certain sense, fields interact only if they are closer than coincident in time. This sounds contradictory, but it demonstrates the difficulty with the position space picture and why it is not expected to be an issue for us. The covariant phase space formalism is primarily concerned with infrared physics, and the nature of interactions at short distances should not matter. The singular nature of SFT interactions in time will be called transgressive locality.
@@ 3.2 Symplectic structure
The boundary conditions (3.16) constrain the form of the sigmoid towards zero momentum as
@@
Let us assume that the sigmoid is a function of time only, $\sigma=\sigma(t)$. We further integrate out spatial coordinates to produce a factor of the spatial volume $V$. Finally, we Fourier transform so that integration over time is replaced with integration over energy. In this way, the leading coefficient of the symplectic form is found to be
@@ 3.3 Energy to leading order $\lambda^{2}$
We calculate the leading coefficient $\Omega_{1}$. This coefficient is somewhat trivial in that it merely determines the energy of a solution to the Klein-Gordon equation. Nevertheless the calculation is instructive as the basic idea generalizes to higher order coefficients.
@@ 3.4 Energy to order $\lambda^{4}$
$K_{3}(0,-i\mu,-i\mu,2i\mu)=-K_{3}(0,i\mu,i\mu,-2i\mu)=-3\Lambda\mu e^{-\frac{3\Lambda\mu^{2}}{2}},$ (3.51c)
$K_{3}(0,-2i\mu,i\mu,i\mu)=-K_{3}(0,2i\mu,-i\mu,-i\mu)=-3\Lambda\mu e^{-\frac{3\Lambda\mu^{2}}{2}},$ (3.51d)
@@ 3.5 Transgressive locality and runaway oscillations
where $\lambda=6\mu^{2}e^{-\mu t_{0}}/g$. To map this solution to the effective field theory we need to express it in momentum space. Then we can apply the stub operator following (3.6) to find
@@
This does not vanish at high energies. In fact, it grows without bound. Therefore, the rolling tachyon does not exist within the space of fields where the action and equations of motion are well-defined.
@@
The solution appears to exist even nonperturbatively because the expansion in powers of $e^{\mu t}$ has infinite radius of convergence. The field slowly falls off the unstable vacuum in the distant past and then transitions to a phase of violent, runaway oscillation. This is shown in figure 3.2. The picture is quite analogous to rolling tachyon solutions found in $p$-adic string theory and open SFT [18, 19, 20]. However, the picture seems incompatible with what we know about the rolling tachyon in scalar field theory. The oscillation continues forever and never settles back to the unstable vacuum.
There is a solution where the oscillation settles to the unstable vacuum. This is simply the time reverse of (3.60). Both solutions have a common origin. They both derive from applying the stub operator to an expansion of the original $\phi^{3}$ solution. The solution (3.60) derives from the expansion in exponential modes, while its time reverse comes from the expansion in inverse exponential modes. In the original $\phi^{3}$ theory these expansions converge to different portions of the same nonperturbative solution. In the effective field theory, however, the expansions have infinite radius of convergence and create different solutions. Another point of view on this comes from the Fourier transform
@@
as shown in figure 3.3. The two contours give respectively the solution (3.60) or its time reverse. It is somewhat puzzling to find two rolling tachyon solutions when there should only be one. We believe that the correct interpretation is that the two solutions are physically equivalent, despite the very different nature of their time evolution. One solution is adapted for describing the early time behavior, and the other the late time behavior.
Figure 3.3: Because the rolling tachyon solution in the effective field theory diverges faster than exponentially for large real energies, to describe the time evolution we must implement a Fourier transform over a contour in the quadrant above or below the real axis. Closing the contours at infinity we obtain an infinite sum of residues at positive or negative integer multiples of $i\mu$. This creates two solutions given respectively as an expansion in powers of $e^{-\mu t}$ or $e^{\mu t}$ with infinite radius of convergence. The former gives a good representation of the evolution of the background towards the future, while the later gives a better representation in the past.
We can speculate whether something analogous is happening in open SFT. The exponential rolling tachyon solution in SFT undergoes violent oscillations towards the future. But a different solution might be needed to reveal the true nature of the decay process at late times. Perhaps the solution ends at the tachyon vacuum, as suggested in [40]. It is interesting to consider the pressure of the rolling tachyon background [16, 41], which is proportional to the function
@@
The exponential rolling tachyon solution in open SFT reconstructs $f(t)$ as an expansion in powers of $e^{t}$, for example though computation of the Ellwood invariant [42, 43]. This expansion however diverges at finite time. To continue further, we may need the other solution to reconstruct the pressure as an expansion in powers of $e^{-t}$. There are indications from [44] that the nature of this solution could be inferred from asymptotic analysis of the rolling tachyon background.
@@ 4.1 Symplectic structure
and the interacting part of the symplectic structure drops out. This simplification is highly dependent on the special geometry of the Witten vertex. There is no clear analogue of Witten’s symplectic structure in closed string field theory, for example. For this reason we are interested in a different construction which extends more readily to other SFTs. We consider the sigmoid
@@
The open string center of mass does not commute through the star product. Therefore the interacting contribution to the symplectic structure will not vanish in our case. However the final results should not depend on how the sigmoid is chosen. Therefore our numbers should agree with those obtained by Cho, Mazel, and Yin based on Witten’s construction. A complication however is that Witten’s symplectic structure is not well-defined. We will elaborate on this in subsection 4.6, and explain how Cho, Mazel, and Yin’s remedy [14] relates to our approach.
@@ 4.4 Oscillator expectation values
This is valid assuming that no divergence appears towards the $y=0$ boundary of moduli space. Divergence will be absent if the propagator acts on states with strictly positive conformal weight. Because $\Psi_{-}^{2}$ carries momentum $2i$, the minimum conformal weight that appears in the star product is
@@
Here we need to be careful with the definition of the propagator because it acts on a state which does not necessarily have positive conformal weight. This is because the momentum cancels between $\Psi_{+}$ and $\Psi_{-}$, rather than adding up as with previous terms. The Schwinger representation may encounter divergences from tachyonic states, where the mode counting operator $N$ vanishes, and from massless states, where the mode counting operator is one. However, massless states will not be present because the rolling tachyon solution is exactly marginal, which has been checked explicitly. Therefore we only need to worry about the tachyon divergence. To deal with this we add and subtract a projector onto the tachyon state:
@@ 4.6 Relation to Cho, Mazel, and Yin
It has often been observed that Witten’s SFT is formally local if the string field is expressed as a function of the position of the string midpoint [59]. Therefore, in principle we can find the symplectic structure in the same way as we would in any local field theory, writing coordinates on a $X^{0}(i,\overline{i})=0$ time slice and computing their canonical momenta from the Lagrangian. The result is the symplectic structure proposed by Witten [15], who wrote it as
@@
The expression clearly fits our general formula (1.1). As mentioned earlier, the interacting contribution drops out because the midpoint coordinate commutes through the star product. Unfortunately, Witten’s symplectic structure is singular. The problem is that the plane wave vertex operators in the sigmoid have nonzero conformal weight. They obey the transformation law
@@
The conformal maps relevant to Witten’s SFT usually have conical singularities at the midpoint. This means that $\partial f(i)$ will be vanishing or infinite. Therefore the symplectic structure is divergent. In the end the difficulty can be traced back to the assumption that Witten’s SFT is local in the midpoint coordinate. Attempts to make this locality manifest encounter serious difficulties from short distance singularities on the worldsheet [60, 61, 62, 63].
@@
The lightlike plane wave vertex operator has vanishing conformal weight and can be inserted at the midpoint. The resulting symplectic structure is
$\Omega=-\frac{1}{2g^{2}}\big\langle\delta\Psi,\gamma^{+}(i,\overline{i})\dot{\sigma}~\delta\Psi\big\rangle,$ (4.117)
where the dot on $\sigma$ indicates the derivative with respect to $X^{+}(i,\overline{i})$. Because the midpoint insertion is finite, the symplectic structure appears to be well-defined. This is related to the fact that Witten’s SFT appears to be local in a more rigorous sense when considering only the lightlike coordinate of the string midpoint. This is the basis of the canonical formulation of Witten’s SFT in midpoint lightcone time [22, 23]. In that context $X^{1}(z,\overline{z})$ is a noncompact free boson subject to Neumann boundary conditions, but for present purposes it is slightly cleaner to assume Dirichlet boundary conditions to avoid issues with momentum conservation and volume divergence in the spatial BCFT. It is interesting to mention that, differently from previous examples, the midpoint-lightcone sigmoid does not act only through multiplication by a function $\sigma(x)$. It mixes fields of different mass level and operates on their Lorentz indices. For this reason, the calculation of the symplectic structure proceeds very differently with the midpoint-lightcone sigmoid than with the center-of-mass sigmoid. Nevertheless the results should be the same. This shows that the choice of sigmoid can be nontrivial and useful, even beyond the simple choice of time slice that we have in more conventional descriptions of the covariant phase space formalism.
@@
Because $\mathcal{O}$ has zero energy the correlation function is nonzero only when the plane wave vertex operator reduces to the identity. Then all contractions between $\mathcal{O}$ and the plane wave vertex operator vanish, and the correlation function effectively factorizes. We assume that the spatial component of $X^{+}$ is subject to Dirichlet boundary conditions, in which case
@@
The final step is to observe that the projection onto the zero energy contribution is no longer necessary because it is enforced automatically by momentum conservation. Then we recover (4.113). The conclusion is that Cho, Mazel, and Yin’s formula can be seen as a reduction of the midpoint-lightcone symplectic form. Independence from the choice of sigmoid then explains why our results must agree with theirs.
We can ask whether our computation of the rolling tachyon energy is more complicated. The midpoint-lightcone formulation has the advantage that interacting contributions to the symplectic structure drop out. However, there are additional contributions from $\Psi_{1}$-$\Psi_{3}$ terms which do not vanish because the sigmoid does not commute through the mode counting operator. Both approaches seem to be viable from a numerical point of view, though our different methods of evaluation make a precise comparison difficult. On the whole our approach compares well to Cho, Mazel, and Yin’s. This is perhaps surprising given the seeming awkwardness of the position zero mode from the point of view of the geometry of the Witten vertex.
The main advantage of the midpoint-lightcone approach is that it is connected to an initial value problem. If the initial value problem is viable, it should give a systematic parametrization of the phase space of open string field theory. Whether it really does remains to be seen. Another path to constructing the phase space is though lightcone gauge, but lightcone gauge in covariant SFT is not completely understood due the appearance of singular Mandelstam diagrams [48, 64]. The explicit construction of phase space in SFT remains a very important open problem.

### v2

# Symplectic structure in open string field theory I: Rolling tachyons

@@ 1 Introduction
In recent work [1] we proposed a formula for the symplectic structure on the phase space of an arbitrary Lagrangian field theory:
@@
The formula assumes that the theory has been expressed in terms of a cyclic $L_{\infty}$ algebra, where $\omega$ is the Batalin-Vilkovisky inner product, $Q_{\Phi}$ is the kinetic operator around a solution $\Phi$, and $\sigma$ is a grade zero operator called the sigmoid. This is the first of three papers [2, 3] whose goal is to test this formula in the context of string field theory (SFT). The hope is that the symplectic structure can serve as a foundation for a Hamiltonian description of the theory, which may give access to new observables in string theory such as energy, topological charge, and black hole entropy. In principle the formula (1.1) applies to both open and closed string field theory. However, we will focus on open (bosonic) strings. Closed SFT is a theory of gravity, which makes extraction of observables (such as energy) more subtle. Some relevant work on diffeomorphism invariance and boundary terms in closed SFT can be found in [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. We leave applications to closed SFT for the future.
An important motivation for our work is the analysis of Cho, Mazel, and Yin [14] who use an older proposal for the symplectic structure due to Witten [15] to compute the energy of rolling tachyon solutions on unstable D-branes [16]. Their results match with boundary state calculations carried out by Kudrna [17]. The agreement, however, requires discarding certain infinities contained within Witten’s formula. In this paper we revisit this calculation using the new symplectic structure. Our ability to confirm these results is a minimum standard for the formalism we are trying to develop.
The divergences in Witten’s symplectic structure stem from the use of the open string midpoint to measure time. In our calculations, we will measure time with the open string center of mass. This is the time coordinate associated to the component fields of the Fock space expansion of the SFT action. With this definition of time, the symplectic form (1.1) is finite. A further benefit is that its evaluation does not rely on special properties of the Witten vertex that have no analogue in other SFTs. A challenge, however, is that Witten’s string field theory is highly nonlocal in the center of mass coordinate. The nature of the nonlocality is also strange. The interactions are in a sense “more than local” in time, as expressed by couplings which grow faster than exponentially at high energies. We refer to this as transgressive locality. A consequence is that fields do not interact in a well-defined way as a function of time. Instead interactions must be expressed in momentum space. The lack of a clear picture of time evolution is worrisome given that we are trying to do Hamiltonian mechanics, but we will see that everything is working.
@@
Finally we consider open string field theory. We compute the symplectic structure for rolling tachyon solutions in Siegel gauge using the center of mass coordinate to measure time. We reduce the expression to first subleading order to a collection of correlation functions on the surface defining the off-shell 4-point amplitude in Siegel gauge. The correlation functions are evaluated using squeezed state oscillator methods developed by Taylor [21], showing good agreement with the results of Cho, Mazel and Yin and very close agreement with boundary state results of Kudrna. The details of our numerical implementation can be found in a Mathematica file accompanying this paper. We conclude by reviewing how the divergences of Witten’s symplectic structure can be eliminated by using the lightcone coordinate of the string midpoint to measure time [22, 23]. Independence from the choice of sigmoid can then be used to explain why our results agree with those of Cho, Mazel, and Yin.
@@ 2 Transgressive locality
The cubic term contains an infinite derivative operator $e^{\frac{\Lambda}{2}\Box}$ which appears (for some $\Lambda$) ubiquitously in interactions between any fields in the Fock space expansion of any covariant string field theory. For this reason nonlocality in string field theory takes a fairly universal form. The infinite derivative operator can be interpreted as a formal power series in $\Box$. However this definition is strictly perturbative in $\alpha^{\prime}$. A more complete definition can be given in terms of a position space integral transform,
@@
The problem with momentum space is that we cannot talk directly about time evolution. This could be worrisome because the consistency of the symplectic structure relies on the assumption that fields propagate and interact in a somewhat localized fashion as a function of time. It might appear that SFT interactions are not like this. However, we will argue that really they are. In fact, in a certain sense, fields interact only if they are closer than coincident in time. This sounds contradictory, but it demonstrates the difficulty with the position space picture and why it is not expected to be an issue for us. The covariant phase space formalism is primarily concerned with infrared physics, and the nature of interactions at short distances should not matter. The singular nature of SFT interactions in time will be called transgressive locality.
@@ 3.2 Symplectic structure
The boundary conditions (3.16) constrain the form of the sigmoid towards zero momentum
@@
Let us assume that the sigmoid is a function of time only $\sigma(t)$. We further integrate out spatial coordinates to produce a factor of the spatial volume $V$. Finally, we Fourier transform so that integration over time is replaced with integration over energy. In this way, the leading coefficient of the symplectic form is found to be
@@ 3.3 Energy to leading order $\lambda^{2}$
We calculate the leading coefficient $\Omega_{1}$. This coefficient is somewhat trivial because it merely determines the energy of a solution to the Klein-Gordon equation. Nevertheless the calculation is useful because it generalizes to higher order coefficients.
@@ 3.4 Energy to order $\lambda^{4}$
$K_{3}(0,-i\mu,-i\mu,2i\mu)=-K_{3}(0,i\mu,i\mu,-2i\mu)=-3\Lambda\mu e^{-3\frac{\Lambda\mu^{2}}{2}},$ (3.51c)
$K_{3}(0,-2i\mu,i\mu,i\mu)=-K_{3}(0,2i\mu,-i\mu,-i\mu)=-3\Lambda\mu e^{-3\frac{\Lambda\mu^{2}}{2}},$ (3.51d)
@@ 3.5 Transgressive locality and runaway oscillations
where $\lambda=6\mu^{2}e^{-\mu t_{0}}/g$. To map this solution to the effective field theory we need to express it in momentum space as $\widehat{\phi}(E)$. Applying the stub operator then gives
@@
This does not vanish at high energies. In fact, it grows without bound. Therefore, the rolling tachyon solution does not exist within the space of fields where the action and equations of motion are well-defined.
@@
The solution appears to exist even nonperturbatively because the expansion in powers of $e^{\mu t}$ has infinite radius of convergence. The field slowly falls off the unstable vacuum in the distant past and then transitions to a phase of violent, runaway oscillation. This is shown in figure 3.2. The picture is quite analogous to rolling tachyon solutions found in $p$-adic string theory and open SFT [18, 19, 20]. However, the physics appears to be incompatible with what we know about the rolling tachyon background in scalar field theory. The oscillations continue forever and never settle back to the unstable vacuum.
There is a solution where the oscillation does settle to the unstable vacuum. This is simply the time reverse of (3.60). Both solutions have a common origin. Both derive from applying the stub operator to an expansion of the original $\phi^{3}$ solution. The solution (3.60) derives from the expansion in exponential modes, while its time reverse comes from the expansion in inverse exponential modes. In the original $\phi^{3}$ theory these expansions converge to different portions of the same nonperturbative solution. In the effective field theory, however, the expansions have infinite radius of convergence and create different solutions. Another point of view comes by inspection of the Fourier transform
@@
as shown in figure 3.3. The two contours give respectively the solution (3.60) or its time reverse. It is strange that the effective field theory has two rolling tachyon solutions when there should only be one. We believe that the correct interpretation is that the two solutions are physically equivalent, despite the very different nature of their time evolution. One solution is adapted for describing the early time behavior, and the other the late time behavior.
Figure 3.3: Because the rolling tachyon solution in the effective field theory diverges faster than exponentially for large real energies, to describe the time evolution we must implement a Fourier transform over a contour in the quadrant above or below the real axis. Closing the contours at infinity we obtain an infinite sum of residues at positive or respectively negative integer multiples of $i\mu$. This creates two solutions given respectively as an expansion in powers of $e^{-\mu t}$ or $e^{\mu t}$ with infinite radius of convergence. The former gives a good representation of the evolution of the background towards the future, while the later gives a better representation in the past.
We can speculate whether something analogous happens in open SFT. The exponential rolling tachyon solution in SFT undergoes violent oscillations towards the future. A different solution might be needed to reveal the nature of the decay process at late times. Perhaps the solution ends at the tachyon vacuum, as suggested in [40]. It is interesting to consider the pressure of the rolling tachyon background [16, 41], which is proportional to the function
@@
The exponential rolling tachyon solution in open SFT reconstructs $f(t)$ as an expansion in powers of $e^{t}$, for example though computation of the Ellwood invariant [42, 43]. This expansion however diverges at finite time. Another solution may be needed to reconstruct the pressure as an expansion in powers of $e^{-t}$. Perhaps the solution can be inferred from asymptotic analysis of the rolling tachyon background, following [44].
@@ 4.1 Symplectic structure
and the interacting part of the symplectic structure drops out. This simplification is highly dependent on the special geometry of the Witten vertex. We are interested in a different construction which extends more readily to other SFTs. We consider the sigmoid
@@
The open string center of mass does not commute through the star product. Therefore the interacting contribution to the symplectic structure will not vanish in our case. This makes our computation of the symplectic structure quite different from Cho, Mazel, and Yin’s, which is based on Witten’s construction. However the final results should not depend on how the sigmoid is chosen. A slight complication is that Witten’s symplectic structure is not really well-defined. We will elaborate on this in subsection 4.6, and explain how Cho, Mazel, and Yin’s remedy [14] relates to our approach.
@@ 4.4 Oscillator expectation values
This is valid assuming that no divergence appears towards the $y=0$ boundary of moduli space. Divergence will be absent if the propagator acts on states with strictly positive conformal weight. Because $\Psi_{-}^{2}$ carries momentum $2i$, the minimum conformal weight which appears in the star product is
@@
Here we need to be careful with the definition of the propagator because it acts on a state which does not necessarily have positive conformal weight. This is because the momentum cancels between $\Psi_{+}$ and $\Psi_{-}$, rather than adding up as with previous terms. The Schwinger representation may encounter divergences from tachyonic states, where the mode counting operator $N$ vanishes, and from massless states, where the mode counting operator is one. However, massless states will not be present because the rolling tachyon solution is exactly marginal. Therefore we only need to worry about the tachyon divergence. To deal with this we add and subtract a projector onto the tachyon state:
@@ 4.5 Results
Cho, Mazel and Yin have demonstrated analytically that the energy computed from the Ellwood invariant must agree to all orders in $\lambda$ with that computed from their version of the symplectic structure. The formal equivalence between their symplectic structure and ours is explained in more detail below. The numerical results demonstrate that these arguments can be trusted.
@@ 4.6 Relation to Cho, Mazel, and Yin
It has often been observed that Witten’s SFT is formally local if the string field is expressed as a function of the position of the string midpoint [59]. Therefore, in principle we can find the symplectic structure in the same way as we would in any local field theory, writing coordinates on an $X^{0}(i,\overline{i})=0$ time slice and computing their canonical momenta from the Lagrangian. The result is the symplectic structure proposed by Witten [15], who wrote it as
@@
The expression clearly fits our general formula (1.1). As mentioned earlier, the interacting contribution drops out because the midpoint coordinate commutes through the star product. Unfortunately, Witten’s symplectic structure is undefined. The problem comes from the fact that the plane wave vertex operators in the sigmoid have nonzero conformal weight. They obey the transformation law
@@
The conformal maps relevant to Witten’s SFT usually have conical singularities at the midpoint. This means that $\partial f(i)$ will be vanishing or infinite, making the symplectic structure undefined. The difficulty can be traced back to the assumption that Witten’s SFT is local in the midpoint coordinate. Attempts to make this locality manifest encounter serious difficulties from short distance singularities on the worldsheet [60, 61, 62, 63].
@@
Because the lightlike plane wave vertex operator has vanishing conformal weight, conformal transformation does not lead to singularity. The symplectic structure appears to be well-defined and takes the form
$\Omega=-\frac{1}{2g^{2}}\big\langle\delta\Psi,\gamma^{+}(i,\overline{i})\dot{\sigma}\,\delta\Psi\big\rangle,$ (4.117)
where the dot on $\sigma$ indicates the derivative with respect to $X^{+}(i,\overline{i})$. The success of this construction is related to the fact that Witten’s SFT is local in a more genuine sense when considering only the lightlike coordinate of the string midpoint. This is the basis of the canonical formulation of Witten’s SFT in midpoint lightcone time [22, 23]. In that context $X^{1}(z,\overline{z})$ is a noncompact free boson subject to Neumann boundary conditions, but we consider Dirichlet boundary conditions so that we do not have to worry about the zero mode. It is interesting to mention that the midpoint-lightcone sigmoid does not act through multiplication by a function $\sigma(x)$. It mixes fields of different mass level and operates on Lorentz indices. This shows how the sigmoid can usefully generalize the concept of time slice.
@@
Because $\mathcal{O}$ has zero energy the correlation function is nonzero only when the plane wave vertex operator reduces to the identity. Then all contractions between $\mathcal{O}$ and the plane wave vertex operator vanish, and the correlation function effectively factorizes. Since we assume that the spatial component of $X^{+}$ is subject to Dirichlet boundary conditions we have
@@
The final step is to observe that the projection onto zero energy is no longer necessary because it is enforced automatically by momentum conservation. Then we recover (4.113). The conclusion is that Cho, Mazel, and Yin’s formula can be seen as a reduction of the midpoint-lightcone symplectic form. Independence from the choice of sigmoid then explains why our results must agree with theirs. Our argument assumes that the worldsheet theory has a spectator free boson $X^{1}$ whose primaries are not excited by the solution. This will not be true, for example, on a ZZ-brane in $c=1$ string theory. It may be possible to strengthen the derivation by considering Witten’s original sigmoid as a limit.
We can ask how our computation of the rolling tachyon energy compares to Cho, Mazel and Yin’s. Their approach has the advantage that interacting contributions to the symplectic structure drop out. However, there are additional contributions from $\Psi_{1}$-$\Psi_{3}$ terms which do not vanish because the sigmoid does not commute through the mode counting operator. From a numerical point of view both approaches seem to be viable, though our different methods of evaluation make a precise comparison difficult. The conclusion is that our calculation of the symplectic structure is quite simple and natural. Perhaps this is a surprise given the seeming awkwardness of the position zero mode from the point of view of the geometry of the Witten vertex. Therefore we do not expect important obstacles to performing similar calculations to other string field theories, most importantly closed SFT.
The main advantage of the midpoint-lightcone formulation is that it is connected to an initial value problem. If the initial value problem is viable, it should give a systematic parametrization of the phase space of open string field theory. Whether it really does this remains to be seen. Another path to constructing the phase space is though lightcone gauge, but this approach is complicated by the appearance of singular Mandelstam diagrams in the process of gauge fixing [48, 64]. The explicit construction of phase space in SFT remains an important open problem.